A 50 year old man has two children. One of the children is 4 years older than the other. Eight years from now, the father will have the same age as the sum of the ages of the children. What is the age of the younger child now?
a) 14 b) 16 c) 19 d) 24 e) 30

Answer by richwmiller(17219)

f=50
x=y+4
58=y+8+y+4+8
58=20+2y
38=2y
19=y
23=x
check
19+8+23+8=58
27+31=58
58=58